the width of rectangle is 5 feet and the diagonal is 9 feet. what is the area of the rectangle?
Answer by texttutoring(324)

This would be easier to solve with a picture, but I don't know how to insert images here. If you want to see an image, send me a message.
You can make a right triangle with the diagonal of the rectangle and its width.
diagonal = 9
width = 5
use Pythagorean theorem to solve for the unknown side, the length:
9^2 = 5^2 + l^2
81-25 = l^2
l= Sqrt(56)
l = 7.48
The area is A=l*w
A= 7.48 * 5
A= 37.4